OpenStudy (anonymous):

\[F = 9\div5C + 32\]

hartnn (hartnn):

good to know :)

OpenStudy (anonymous):

Can you solve for C using the formula?

hartnn (hartnn):

ite Centigrade to faranheit conversion

hartnn (hartnn):

sure, i can, can u ?

OpenStudy (anonymous):

No thats why I asked, I need help

hartnn (hartnn):

subtract 32 from both sides, what u get ?

OpenStudy (anonymous):

\[F- 32 = 9\div5C\]

hartnn (hartnn):

right, now can u flip the fraction? like if a/b=c/d then b/a=d/c

OpenStudy (anonymous):

\[1\div F-32 = 5\div 9C \] is that correct?

hartnn (hartnn):

just don't forget to keep brackets 1/(F-32)=5C/9

OpenStudy (anonymous):

Okay, do I cross multiple now?

hartnn (hartnn):

now you can multiply both sides by 9/5 so that you isolate C on right side

OpenStudy (anonymous):

when I multiply 9/5 by both sides, do I cross multiple or multiply numerator by numerator and denominator by denominator?

OpenStudy (anonymous):

multiply*

hartnn (hartnn):

(9/5)(1/F-32) = 5C/9 * 9/5

OpenStudy (anonymous):

so whats the final answer?

hartnn (hartnn):

on right side, u get only C on left u get \(\large \frac{9}{5(F-32)}\)

hartnn (hartnn):

thats your final answer

OpenStudy (anonymous):

Thanks so much for your assistance !

hartnn (hartnn):

welcome :)

hartnn (hartnn):

can i clarify something?